The probability that a trainee will remain with a company is 0.6. The probability that an employee earns more than Rs.10, 000 per year is 0.5. The probability that an employee is a trainee who remained with the company or who earns more than Rs.10, 000 per year is 0.7. What is the probability that an employee earns more than Rs.10, 000 per year given that he is a trainee who stayed with the company.
Answer by stanbon(48568)

The probability that a trainee will remain with a company is 0.6.
P(R) = 0.6
-----------------
The probability that an employee earns more than Rs.10, 000 per year is 0.5
P(10K) = 0.5
-----------------
The probability that an employee is a trainee who remained with the company or who earns more than Rs.10, 000 per year is 0.7
P(R or 10K) = 0.7
----------------------------
What is the probability that an employee earns more than Rs.10, 000 per year given that he is a trainee who stayed with the company.
P(10K | R) = P(10K and R)/P(R)
---
Note: P(10K and R) = P(R)+P(10K)-P(10K or R) = 0.6+0.5-0.7 = 0.4
----------
Therefore P(10K |R) = 0.4/0.6 = 2/3
========================================
Cheers,
Stan H.